October 17, 2013 FDC meeting
- Installation preparations  (all)
- Engineering (Bill)
- Electronics (Chris, Nick)
- Tests in EEL126 FDC E-log (Beni, Lubomir)
- New F1TDC
- Results from the full electronics test
Participants: Eugene, Bill,Dave, Chris, Simon, Fernando, Beni, and Lubomir.
- Current status of the FDC related stuff in the Hall shown with pictures linked above. Lubomir and Dave discussed with Tom today the work schedule in the hall. As of today, the plan is to move the FDC to the hall next week, earliest on Tuesday (Oct. 22). By the end of the day they will finish with BCAL cabling and the cooling rings will be installed tomorrow. On Monday all the platform will be cleaned and the survey people will be aligning the top and bottom rails on the first cart and inside BCAL. It turned out the top part of the cart had to be shifted to match the top rail inside BCAL, but the side rails have not been touched. Additional plates on the bottom of all the carts had to be installed to prevent the carts from tilting. By some reasons the top and bottom rails inside BCAL have been replaced by longer once reaching the downstream end of the magnet, but they are not needed that long? (Bill will check this).
- The gas system panel will be installed on the platform at the right side of the magnet at the downstream end. 1/2" tube will go to the FDC and CDC from there. The gas controls will be in a rack at the floor at the left-upstream side of the magnet (to avoid noise in the electronics at the right racks). The cooling system will be at the downstream side of the magnet on the floor. We discussed if the cooling system and the solenoid valves in the gas panel will operate without problems in ~50 gauss magnetic field. Bill will look at this and if possible wants to test them with a field. At the beginning we will use Nitrogen from a dewar to flush the chambers, but by November 11th we want to have at least a temporary gas system with Ar and CO2 to start testing with HV. Casey is working on the chilled water piping needed by November 4th when we will start testing the LV cables.
- As of now the plan is to move FDC middle of next week. Then (for one week at least) there will be a survey and also dry runs of moving FDC in and out of the magnet to make sure the rails are OK and it returns to the same position. Approximately on November 4th we will start connecting and testing the LV cables. Fernando: by that time all the LV, HV and TDC crates will be installed in the racks. One week later, Nov.11, we plan to start connecting the signal cables. At that time we will need the mini DAQ to test the cable connections and Beni is working on this. The mini-DAQ will be placed at the back wall near the pair-spectrometer. For that Beni needs a crate with fADC and TDC with self-triggering feature. Fernando: it's possible but right now Cody has the full crate testing as a priority.
- The mesh cylinder has not been assembled yet and Tom is looking for the parts. Bill will take care of this.
- Nick needs the last signal cables from 126 that has not been labeled yet. Most of them are packed in boxes and ready to be moved but 35 are used by Beni in the test F1TDC test, so these will be taken next week. Vlad installed the last pre-amp on the FDC that had to be replaced.
Tests at EEL126
- Beni is testing the spare cells with the new F1TDCs and will be ready by Monday.
- Beni showed results of deviation of the cathode reconstructed wires from the nominal position as a function of the wire number (Elog entry 118) for the test with full electronics. He is fitting these with quadratic polynomial to find correction to the strip rotations and offsets and soon will have these results.
- Lubomir was also looking at the same deviations from the nominal wire positions, but as a function of both x and y coordinate (y along the wires). These deviation plots (entry 119) were fitted with planes, each having three parameters. Thus for each cell we can use these three parameter to find corrections to the geometry: in this case two rotations, of the top and bottom strips, and one translation of both strip planes in direction perpendicular to the wires. After applying these corrections (see table in Entry 120) the deviations from the nominal are much smaller.
- Time to distance relation was estimated from the cathode information only (Entry 121). For a straight track through the package, one can use the charge positions on the cell 1 and 4 (with parallel wires and strips) to project this on a middle cell (2 or 3) and the crossing with the similar line from the charge on this cell will give us the hit position. The distance of the hit to the wire is then related to the drift time. At zero drift time the distance is not zero, which can be corrected by a shift of the strips along the wires by 2.0/1.5 mm for cell 2/3. In all these corrections the cathode foil elongation is not considered but is effectively taken into account by the rotations. To separate these effects it will be very important to measure the cathode foil deformation; we have one production cathode that is not used in the packages. Bill wants to do this on the stringing table using the step motor, when the controller is back from the tagger magnet measurements.